THis is simple and I can ge tthe answer but not the formula. Ed has $1.15. He has 43 more pennies than nickels. How many pennies and nickels does he have. THe answer is 12 nickels and 55 pennies but i just randomly guessed. It has to be easier than that
Answer by scott8148(6628)

"He has 43 more pennies than nickels" ___ p = n + 43

"Ed has $1.15" ___ 5n + p = 115 ___ nickels worth 5 , pennies worth 1

substituting for p ___ 5n + (n + 43) = 115 ___ 6n + 43 = 115 ___ 6n = 72 ___ n = 12

substituting ___ p = (12) + 43 = 55